[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon Jul 17, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Mon Jul 17 13:40:20 UTC 2017


Here is today's summary of NHC computer operations:
--- Mello

----------

TAFB:
1. The Euro wave model is now updating un GFE.
2. SST data is missing from GFE. Mello will investigate.
3. The collaboration sever in AWIPS2 is stable after they switched us to a  
new Server which seems to have solved the problem. Mello will follow up  
this week.
HSU:
1. The satellite data is flowing again into NAWIPS. Directories are being  
monitored by TSB and will be manually purged if needed. However, TSB will  
look into fixing the purge scripts this week.

2. The problems with Ridges on Sunday appear to be related to the filled  
satellite partition. There have been no problems since Sunday after we  
fixed the data flow issues.
Other items:

NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170717/16966a54/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list